OpenStudy (anonymous):

find the measure of an interior angle and the measure of an exterior angle for the regular polygon. 20-gon.

OpenStudy (anonymous):

i guess jimmyrep solved this for u. but yea, it is 160 and 20 , exterior and interior respectively :)

OpenStudy (amistre64):

180 +(360/20) = exterior angles = 198 360-198 = interior angles = 162

OpenStudy (amistre64):

that last bit was for an 18-gon

OpenStudy (amistre64):

180 +(360/18) = 200 ext 360 - 200 = 160 ints